OpenStudy (anonymous):

Can someone get me started on this? I don't want the whole answer I just want the first step. Solve 43a + 10 - 26a = 27. Do you add 26a to both sides or subtract 43a or what?

OpenStudy (anonymous):

43a+10-26a=27 17a+10=27 17a=17 a=1 Do you understand from the steps I wrote out?

OpenStudy (anonymous):

How did you get the 17a?

OpenStudy (anonymous):

You want to combine like terms. You have 2 like terms. You have #a & #, so you put them together. 43a-26a=17a. Thats how you get 17a+10=27. Then combine the #s. 27-10=17 (You subtract because you are moving the 10 from one side to the other so you change signs, in this case +10 to -10. Thats how you get 17a=17. Then because 17 is with a it is saying 17*a. Therefore you do the opposite to the other side. 17/a=17/17. THat leaves you with a=1.